Math 550 (Section 1)  University of South Carolina
D. Meade  Fall 1997

Homework Problems for Chapters 1 and 2 Review of Math 241

Section Page Problems
1.1  22 # 12, 14^{∗}, 19, 23
1.2  36 # 5, 14^{∗}, 15, 19, 20^{∗}
1.3  53 # 5, 7, 10, 15, 24^{∗}, 25, 26
1.4  63 # 4, 8^{∗}, 9a, 11
Review 75 # 15, 16, 25, 26, 34^{∗}, 42^{∗}, 43, 44
2.1  91 # 5, 6^{∗}, 7, 8
2.2  110 # 5, 6, 15^{∗}
2.3  122 # 1, 5, 8, 10^{∗}
2.4  132 # 1, 2, 3, 8, 11^{∗}, 18
2.5  141 # 3, 8^{∗}, 12, 17ac
2.6  152 # 2, 5, 6, 15, 22^{∗}
Review 164 # 8^{∗}, 10^{∗}, 22, 40, 47^{∗}

\underline{Notes}

• Problems marked with an asterisk (^{∗}) will be graded. The due date will be announced in class.

• These problems have been selected as the minimum amount of work that you should be doing to prepare for the remainder of this course. Some problems will be straightforward computations similar to ones that you did in Math 241; others might be a little different – but should not be difficult once you re-read the text.

• You will need to read the text. The lectures will provide the information needed to complete some problems, but you may need to read the text (or other reference material) in order to solve some of the problems.

• If you need more practice, there are many other problems in the text that are similar to the ones I have assigned.

• Be sure you know how to do all of these problems. Do not hesitate to ask questions – in class, by e-mail, or during office hours.